# CineTrans: Learning to Generate Videos with Cinematic Transitions via Masked Diffusion Models

### v1

@@ E.1 Prompt Details
In Section 6.2, we present a set of prompts generated by GPT-4o (Achiam et al., 2023) to evaluate the performance of multi-shot video generation. Figure 18(a) provides the prompt used to guide GPT.
@@ E.2 Metric details
Transition Control. For transition control, we define the Transition Control Score to measure whether the shot count in the generated video aligns with the specified. Specifically, the generated video is first processed by Transnetv2 for shot segmentation, yielding the number of shots, denoted as $s_{\mathrm{generated}}$, while the prompt specifies a target shot count, $s_{\mathrm{specified}}$. The Transition Control Score can be computed according to Equation 6.
$x=\dfrac{s_{\mathrm{generated}}}{s_{\mathrm{specified}}}$ (5)
$\mathrm{Transition\,Control\,Score}=\dfrac{x^{k}}{e^{k(x-1)}}$ (6)
Figure 19 visualizes the calculation method of Equation 6 for the Transition Control subpanel. In practice, when $x<1$, $k$ is set to 2, and when $x\geq 1$, $k$ is set to 1.6. Given that the prompts used for evaluation all specify multiple transitions, the score is set to 0 when the generated video consists of a single shot. When the shot count in the generated video matches the specified value, the score is recorded as 1. More generally, the score is determined based on the absolute difference from the specified value.
Temporal consistency. In terms of temporal consistency, we consider both intra-shot consistency and inter-shot consistency. Intra-shot consistency treats each shot as a separate video and calculates the metric between adjacent frames using a method similar to that in VBench (Huang et al., 2024). The focus of this section is on inter-shot consistency. As for frame extraction, we use the middle frame of each shot for calculation. However, if the video does not generate multiple shots, inter-shot consistency cannot be evaluated. As shown in Table 2, the original LaVie (Wang et al., 2024) lacks the ability to generate transitions, and therefore its inter-shot consistency metric does not have a corresponding value.
@@
$M=\frac{1}{2}(P+Q)$ (7)
$\mathrm{JS}(P\,\|\,Q)=\dfrac{1}{2}D(P\,\|\,M)+\dfrac{1}{2}D(Q\,\|\,M)$ (8)
$\mathrm{JSD}=\sqrt{\mathrm{JS}(P\|Q)}$ (9)
@@ F.1 Failure Case
Figure 20: Failure case with low consistency, which probably results from insufficient training.
The role of the mask mechanism in controlling the occurrence of transitions is relatively stable. Even in cases where the generated videos does not exactly match the expected shot count, this can be attributed to relatively small compositional differences between shots, which, despite being perceptible to the human eye as content jumps, cannot be detected by the shot segmentation model. However, CineTrans still exhibits failure cases with low consistency, as illustrated in Figure 20, which may be due to insufficient training or the need for more detailed character ID annotations in the dataset.

### v2

@@ E.1 Prompt Details
In Section 5.2, we present a set of prompts generated by GPT-4o (achiam2023gpt) to evaluate the performance of multi-shot video generation. Figure 26(a) provides the prompt used to guide GPT.
@@ E.2 Metric details
Transition Control. For transition control, we define the Transition Control Score to measure whether the shot count in the generated video aligns with the specified count, as formulated in Equation 5.
Figure 27 visualizes the calculation method for the Transition Control subpanel. In practice, when $x<1$, $k$ is set to 2, and when $x\geq 1$, $k$ is set to 1.6. Given that the prompts used for evaluation all specify multiple transitions, the score is set to 0 when the generated video consists of a single shot. When the shot count in the generated video matches the specified value, the score is recorded as 1. More generally, the score is determined based on the absolute difference from the specified value.
Temporal consistency. In terms of temporal consistency, we consider both intra-shot consistency and inter-shot consistency. Intra-shot consistency treats each shot as a separate video and calculates the metric between adjacent frames using a method similar to that in VBench (huang2024vbench). The focus of this section is on inter-shot consistency. As for frame extraction, we use the middle frame of each shot for calculation. However, if the video does not generate multiple shots, inter-shot consistency cannot be evaluated. As shown in Table 1, the original LaVie (wang2024lavie) lacks the ability to generate transitions, and therefore its inter-shot consistency metric does not have a corresponding value.
@@ F.1 Failure Case
Figure 28: Failure case with similar composition consistency, which probably results from insufficient training.
The role of the mask mechanism in controlling the occurrence of transitions remains relatively stable. Nevertheless, in occasional cases (particularly under the training-free setting), the compositions of different shots may become overly similar, as illustrated in Figure 28. Although such content changes are perceptible to the human eye, the shot segmentation model does not recognize them as multi-shot videos due to the high compositional similarity, and the resulting visual experience also deviates from the film editing style. These cases can be attributed to insufficient training or the absence of such data types in the training set, which prevents the model from generating videos of the same scene with varied compositions. Potential improvements may come from expanding the dataset or further training.
@@ F.2 Future Work
  • Fine-grained consistency presents an area for further improvement. While our method demonstrates strong overall consistency, achieving fine-grained alignment across shots remains challenging due to the intricate spatial understanding required. This limitation is primarily due to the lack of detailed background and scene context annotations in the training data. Future work will focus on incorporating more granular annotations and leveraging advanced model designs to address this challenge.
Appendix G Use of Large Language Models
We clarify the involvement of large language models (LLMs) in the preparation of this work. LLMs are not used for research design, methodological decisions, or experimental analysis. Their use is limited to two aspects: (i) improving the clarity and readability of the manuscript through language editing, and (ii) generating evaluation prompts during the assessment phase, as explicitly noted in Section 5.2. All substantive research contributions, including the conception of the problem, model design, implementation, and analysis, are conducted entirely by the authors.